# TF-SASM: Training-free Spatial-aware Sparse Memory for Multi-object Tracking

### v1

@@ IV-D1 Effectiveness of each module
To initiate our experimentation, we systematically incorporated each module into the baseline model, MOTRv2 [motrv2], in an iterative manner, leading to the results depicted in Table II. Our analysis indicates that integrating the Spatial-aware Sparse Memory (SASM), as elaborated in Section III-C, improves the IDF1 score by 1.0 and the AssA score by 1.0. Nevertheless, we observed performance degradation in specific videos due to noise stemming from object feature overlap within the stored memory. We introduce the Overlapping-aware Feature Selector (OFS) detailed in Section III-D to tackle this issue. This addition substantially improves tracking consistency by enhancing both detection accuracy (DetA by 0.3) and tracking accuracy (AssA by 0.4).
@@ IV-D2 Alternative design comparison
Our work introduces two memory modules: the Spatial-aware Sparse Memory (SASM) and the Overlapping-aware Feature Selector (OFS). We conducted a series of empirical studies to optimize each module’s design. The first investigation focused on memory density, comparing feature storage at every frame (dense memory) with a sparse approach (Table III). In both experiments, the maximum stored features were capped at 10. In particular, dense memory exhibited significantly lower performance across all metrics in the DanceTrack validation set. We attribute this disparity to rapid object deformation. In particular, for stationary objects, dense memory features become biased towards the current state, leading to missed detections when objects undergo sudden deformations, a frequent occurrence in the DanceTrack dataset.
TABLE III: Comparison between Dense and Sparse memory
@@
Sparse 66.1 (+1.6) 79.1 (+0.8) 55.3 (+2.0) 69.4 (+2.2)
Our next investigation addressed the challenge of handling overlapping object features during storage. As described in Section III-D, two approaches were considered: delaying storage until a frame with a good location of the object is encountered or storing the feature at the best location since the last storage. Our findings revealed that the delay approach introduced a significant drawback: large gaps between feature storage events due to frequent object overlaps (Table IV). Both experiments utilized the same SASM module to ensure a fair comparison.
TABLE IV: Comparison between Delaying and OFS strategy
@@
OFS 66.5 (+0.5) 79.4 (+0.4) 55.8 (+0.5) 69.5 (+0.7)
Figure 3: Qualitative results for tracking consistency
Figure 4: Qualitative results for tracking re-id performance
@@ IV-D3 Hyperparameters
We conducted thorough experiments to explore our algorithm’s sensitivity to hyperparameters. Initially, we focused on determining the optimal number of frames to store in memory for each object. Our investigation revealed that utilizing 10 frames yielded the most favorable outcomes. Conversely, as detailed in Table V, alternative configurations resulted in notably inferior performance. This discrepancy arises from the inherent challenge of aligning memory features with object features, a task complicated by the absence of explicit training in our algorithm. The more extended number of frames ($\geq 15$ frames) of memory exacerbates this issue by introducing sparse data and noise from tracking failures, leading to a substantial mismatch between memory and object features. Conversely, a paucity of frames (5 frames) constrains the algorithm’s capacity to capture sufficient temporal information.
TABLE V: Comparison between difference memory length
# frames HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
5 66.0 79.1 55.2 68.9
10 66.5 79.4 55.8 69.5
15 65.8 79.1 54.9 68.5
20 65.6 78.6 54.9 68.6
Furthermore, we conducted experiments to evaluate the impact of adjusting the threshold on the algorithm’s performance as in Table VI. Our analysis indicates that setting the threshold ($\epsilon$) to 0.05 closely aligns the algorithm’s performance with that of Dense memory, as illustrated in Table III. However, surpassing a threshold of 0.1 relative to the image size leads to excessively prolonged movement distances. This delay in storing image features in memory exacerbates the disparity between stored and current object features, resulting in a rapid performance decline hovering around 65.8%. Nevertheless, even with highly sparse memory, the algorithm exhibits performance enhancements compared to the baseline, achieving 65.4%.
TABLE VI: Comparison between difference moving threshold
$\epsilon$ HOTA ($\uparrow$) DetA ($\uparrow$) AssA ($\uparrow$) IDF1 ($\uparrow$)
@@ IV-E Qualitative result
Furthermore, in Figure 4, observe the individual with ID 4, who occasionally experiences mis-detection in varying light conditions. Despite this challenge, our methods maintain the object’s ID even after the model redetects the object in the final frame. In contrast, MOTRv2 not only struggles to associate the misdetected object but also erroneously links the person with ID 3 to the one with ID 4, creating a new ID for the individual previously identified as ID 4 in the preceding frame.
In conclusion, our qualitative analysis highlights the robustness and efficacy of our method in tracking objects in challenging scenarios. From maintaining consistency during viewpoint changes to preserving object IDs after missed detections, our approach demonstrates superior performance to existing models like MOTRv2.

### v2

@@ IV-D1 Effectiveness of each module
To initiate our experimentation, we systematically incorporated each module into the baseline model, MOTRv2 [motrv2], in an iterative manner, leading to the results depicted in Table II. Our analysis indicates that integrating the Spatial-aware Sparse Memory (SASM), as elaborated in Section III-B, improves the IDF1 score by 1.0 and the AssA score by 1.0. Nevertheless, we observed performance degradation in specific videos due to noise stemming from object feature overlap within the stored memory. To tackle this issue, we introduce the Overlapping-aware Feature Selector (OFS) detailed in Section III-C. This addition substantially improves tracking consistency by enhancing both detection accuracy (DetA by 0.3) and tracking accuracy (AssA by 0.4).
@@ IV-D2 Alternative design comparison
Our work introduces two memory modules: the Spatial-aware Sparse Memory (SASM) and the Overlapping-aware Feature Selector (OFS). We conducted a series of empirical studies to optimize each module’s design. The first investigation focused on memory density, comparing feature storage at every frame (dense memory) with a sparse approach (Table III). In both experiments, the maximum stored features were capped at 10. In particular, dense memory exhibited significantly lower performance across all metrics in the DanceTrack validation set. We attribute this disparity to rapid object deformation. For stationary objects, dense memory features tend to bias towards the current state, often missing sudden deformations common in the DanceTrack dataset. In contrast, our method stores fewer features during minor object deformations while preserving longer temporal information, aiding in more accurate object state associations.
Our next investigation focused on addressing the challenge of handling overlapping object features during storage. As described in Section III-C, we explored two approaches: delaying storage until encountering a frame with optimal object positioning or storing the feature at the best location since the last storage. Our findings uncovered a significant drawback with the delay approach: large gaps between feature storage events due to frequent object overlaps (Table III). Both experiments used the SASM module to ensure a fair comparison.
TABLE III: Comparison between Memory Types and Different Strategies
Design Performance Metrics ($\uparrow$)
HOTA DetA AssA IDF1
@@
Sparse (Ours) 66.1 (+1.6) 79.1 (+0.8) 55.4 (+2.1) 69.4 (+2.2)
@@
OFS (Ours) 66.5 (+0.5) 79.4 (+0.4) 55.8 (+0.5) 69.5 (+0.7)
@@ IV-D3 Hyperparameters
We conducted thorough experiments to explore our algorithm’s sensitivity to hyperparameters. Initially, we focused on determining the optimal number of frames to store in memory for each object. Our investigation revealed that utilizing 10 frames yielded the most favorable outcomes. Conversely, as detailed in Table IV, alternative configurations resulted in notably inferior performance. This discrepancy arises from the inherent challenge of aligning memory features with object features, a task complicated by the absence of explicit training in our algorithm. The more extended number of frames ($\geq 15$ frames) of memory exacerbates this issue by introducing sparse data and noise from tracking failures, leading to a substantial mismatch between memory and object features. Conversely, a paucity of frames (5 frames) constrains the algorithm’s capacity to capture sufficient temporal information.
We also evaluated the impact of adjusting the threshold on our algorithm’s performance (Table IV). Setting the threshold ($\epsilon$) to 0.05 aligns closely aligns the algorithm’s performance with the baseline. However, thresholds exceeding 0.1 relative to image size result in prolonged movement distances, delaying feature storage in memory. This exacerbates disparities between stored and current object features, leading to a performance decline to around 65.8%. Nonetheless, even with highly sparse memory, our algorithm shows performance improvements compared to the baseline, achieving 65.4%.
TABLE IV: Comparison between different parameters: moving threshold vs memory length
Parameter Performance Metrics ($\uparrow$)
HOTA DetA AssA IDF1
Moving Threshold ($\epsilon$)
@@ IV-E Qualitative result
Figure 3: Qualitative results for tracking consistency
@@ V Conclusion
Acknowledgements
This research is funded by University of Science, VNU-HCM under grant number T2023-93. Furthermore, Nguyen Quang Thuc was also funded by the Master, PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2022.ThS.JVN.10.